# Tighter Privacy Auditing of DP-SGD in the Hidden State Threat Model

### v2

@@ Abstract
Machine learning models can be trained with formal privacy guarantees via differentially private optimizers such as DP-SGD. In this work, we focus on a threat model where the adversary has access only to the final model, with no visibility into intermediate updates. In the literature, this “hidden state” threat model exhibits a significant gap between the lower bound from empirical privacy auditing and the theoretical upper bound provided by privacy accounting. To challenge this gap, we propose to audit this threat model with adversaries that craft a gradient sequence designed to maximize the privacy loss of the final model without relying on intermediate updates. Our experiments show that this approach consistently outperforms previous attempts at auditing the hidden state model. Furthermore, our results advance the understanding of achievable privacy guarantees within this threat model. Specifically, when the crafted gradient is inserted at every optimization step, we show that concealing the intermediate model updates in DP-SGD does not amplify privacy. The situation is more complex when the crafted gradient is not inserted at every step: our auditing lower bound matches the privacy upper bound only for an adversarially-chosen loss landscape and a sufficiently large batch size. This suggests that existing privacy upper bounds can be improved in certain regimes.
@@ 1 Introduction
In this work, we consider a natural relaxation where intermediate models are concealed and only the final model is released. This threat model, often referred to as hidden state, is particularly relevant in practice, encompassing scenarios such as open-sourcing a trained model by publishing its weights. From a theoretical perspective, recent work has demonstrated that the hidden state model can yield significantly improved privacy upper bounds for DP-SGD compared to those derived through standard composition (Ye and Shokri, 2022; Altschuler and Talwar, 2022). This improvement can be attributed to the phenomenon of privacy amplification by iteration (Feldman et al., 2018; Balle et al., 2019): in a nutshell, the privacy of a data point used at earlier stages of the optimization process improves as subsequent steps are performed. However, these results only hold for convex problems, leaving a pivotal question unanswered: Does the privacy of non-convex machine learning problems improve when intermediate models are withheld? While empirical lower bounds obtained through privacy auditing (Nasr et al., 2021; 2023; Steinke et al., 2023) suggest that it might be the case, it remains unclear whether the gap with upper bounds is because genuine privacy amplification occurs or because the existing adversaries are suboptimal, leading to loose privacy lower bounds. This raises another crucial question: How can one design worst-case adversaries when the intermediate models are concealed?
@@
Our contributions. Leveraging the observation that all privacy accounting techniques for DP-SGD protect against worst-case gradients, we adapt the idea of gradient-crafting adversaries (Nasr et al., 2021) to the hidden state model. Instead of crafting a data point (canary) that gets added to the training set as in prior attempts to audit the hidden state model (Nasr et al., 2023; Steinke et al., 2023), our adversaries craft a sequence of gradients prior to the execution of the algorithm (i.e., without knowledge of the intermediate models). The crafted gradients are then added to gradients computed on real training points to yield the highest possible privacy loss for the final model. In other words, our adversaries abstract away the canary and directly decide the gradient it would have produced when inserted at a given step of DP-SGD. But how do we pick the gradient sequence leading to the worst-case leakage?
@@
In the scenario where crafted gradients are inserted at every iteration of DP-SGD, we demonstrate that gradient-crafting adversaries which allocate the maximum magnitude permitted by DP-SGD to a single gradient dimension are optimal: they imply privacy lower bounds that match the known upper bounds given by numerical composition (Gopi et al., 2021). Therefore, our results reveal that releasing only the final model does not amplify privacy in this regime. In the case of small models, we achieve these tight privacy auditing results by carefully selecting the gradient dimension, whereas for over-parameterized models, the dimension selection can be arbitrary.
@@
When the crafted gradient is not inserted at every step, we find that the above adversaries still outperform canary-crafting adversaries by a significant margin but cannot reach the privacy upper bounds. We show that part of the gap can be attributed to the inability of known adversaries to influence the gradients on real training data in steps where the crafted gradient is not inserted. To address this gap, we design an adversary that crafts both the gradient and the loss landscape. Our results uncover two distinct regimes: (i) when the batch size is large compared to the noise variance, our adversary proves to be optimal, yielding a privacy lower bound that matches the known upper bound and thus demonstrating the absence of privacy amplification; and (ii) when the batch is small relative to the noise variance, we show strong evidence of privacy amplification for non-convex problems, although the effect is qualitatively weaker than in the convex case. Our findings advance the understanding of privacy guarantees in the hidden state model, and lay the groundwork for the design of better privacy accounting techniques for this threat model.
@@ 2.1 Differential Privacy and DP-SGD
The workhorse of private machine learning is the Differentially Privacy Stochastic Gradient Descent (DP-SGD) algorithm (Song et al., 2013; Bassily et al., 2014; Abadi et al., 2016). Let $D$ be the training dataset, $\theta$ the model parameters and consider the standard empirical risk minimization objective $\min_{\theta}\frac{1}{|D|}\sum_{x\in D}\ell(\theta;x)$ where $\ell$ is a loss function differentiable in its first parameter. DP-SGD follows similar steps as standard (non-private) SGD but ensures differential privacy by (i) bounding the contribution of each data point to the gradient using clipping and (ii) adding Gaussian noise to the clipped gradients.
Formally, starting from some initialization $\theta_{0}$, DP-SGD performs $T$ iterative updates of the form:
@@ 2.2 Threat Models for DP-SGD
DP protects against an adversary that observes the output of either $\mathcal{M}(D)$ or $\mathcal{M}(D^{\prime})$ and seeks to predict whether the dataset was $D$ or $D^{\prime}=D\cup\{x\}$, i.e., whether $x$ was included in the input dataset (a.k.a. membership inference). In this context, the threat model specifies which information is observable/known by the adversary. For DP-SGD, in addition to the final model $\theta_{T}$, threat models typically consider that the adversary has access to (and potentially controls) the model architecture, the loss $\ell$, the initialization $\theta_{0}$, the dataset (up to the presence of $x$) and differ in which internal states of DP-SGD are observable by the adversary. Below, we recall the two threat models relevant to our work.
Standard threat model. Standard privacy accounting techniques analyze DP-SGD as a composition of (potentially subsampled) Gaussian mechanisms (Abadi et al., 2016; Gopi et al., 2021; Doroshenko et al., 2022). Composition allows the adversary to observe all intermediate models $\theta_{1},...,\theta_{T-1}$ (in addition to $\theta_{0}$ and $\theta_{T}$). Previous work has shown that existing privacy accounting techniques are tight in this threat model (Nasr et al., 2021; 2023). However, revealing intermediate models is unnecessary in most realistic deployment settings and may hurt the privacy-utility trade-off, motivating the study of the hidden state threat model.
@@ 2.3 Auditing Differential Privacy
Privacy accounting only provides upper bounds on the $(\varepsilon,\delta)$-DP parameters, and these bounds are not always tight. Leveraging the relation between DP and the performance of membership inference attacks, privacy auditing aims to produce lower bounds on the DP parameters by instantiating concrete adversaries in various threat models (Jagielski et al., 2020; Nasr et al., 2021). When these lower bounds match the upper bounds given by privacy accounting, we can conclude that the privacy analysis is tight; when they do not, it is possible to improve the privacy accounting and/or the attacks. A privacy auditing pipeline can be broken down into two components: an adversary and an auditing scheme. The high-level process is shown in Algorithm 1.
@@
     $b_{i}\leftarrow\text{Ber}(1/2)$draw a random bit}
@@ 3 Related Work in Differential Privacy Auditing
Prior attempts at auditing the hidden state model used adversaries employing a loss-based attack that targets a canary obtained by flipping the label of a genuine data point (Nasr et al., 2021; 2023; Steinke et al., 2023; Annamalai and Cristofaro, 2024). The idea is to generate an outlier from an in-distribution data point so that the loss of the model on this canary is high when it is not part of the training set but drops significantly when used during training. Auditing the hidden state model with these adversaries yields privacy lower bounds that exhibit a significant gap compared to privacy accounting upper bounds for the standard threat model (Nasr et al., 2021; 2023; Steinke et al., 2023; Annamalai and Cristofaro, 2024). This gap has two possible explanations: (i) these adversaries are suboptimal and stronger adversaries exist in the hidden state model; and/or (ii) the upper bounds are not tight. Understanding the reasons for this gap, so as to gain better knowledge of the properties of DP-SGD in the hidden state model, is the main motivation for our work. We demonstrate that gradient-crafting adversaries can provide tighter auditing results, thereby proving the validity of hypothesis (i). We show that in certain scenarios, our auditing results match the privacy upper bounds, indicating that releasing intermediate models does not increase the privacy loss. However, we also identify regimes where a gap remains, thus providing compelling evidence in support of hypothesis (ii). Specifically, our findings suggest the presence of a privacy amplification by iteration phenomenon in the non-convex setting, albeit weaker compared to what is theoretically established for the convex case (Feldman et al., 2018; Balle et al., 2019), but less restrictive than the results of Asoodeh and Diaz (2023). The latter require projections onto a convex set of bounded diameter or a strong enough decay term on the parameters, which are unrealistic assumptions in practical scenarios involving non-convex and over-parameterized models like the ones considered in our work.
@@ 4 Gradient-Crafting Adversaries for the Hidden State Model
In all threat models, DP-SGD enjoys a simplified interpretation as a sequence of $T$ sum queries with bounded terms, where the $t$-th query corresponds to summing the clipped gradients over the mini-batch $B_{t}$ in Equation 2. Therefore, privacy accounting techniques for DP-SGD do not have the construct of a fixed set of data points that generate gradients but instead account for any possible gradients with bounded norm $C$. This highlights a key limitation of adversaries used so far in the hidden state model: due to the non-convexity of the objective function, any sequence of gradients for the canary $x^{*}$ is likely to be possible, but how to craft a worst-case $x^{*}$ is unclear (and in fact out of scope for privacy auditing). Instead, we allow the adversary to directly craft any gradient sequence, considering that there could be a data point that would generate that sequence.
@@
  • Saturating the gradient norm: The canary point crafted by prior adversaries is not guaranteed to saturate the gradient clipping threshold throughout training. The adversary’s performance becomes architecture-dependent: for example, a small convolutional neural network has a small canary norm at the start of the training compared to a ResNet, as observed in Figure 1. Thus, for a tight audit, one must tune the clipping threshold according to the canary gradient norm, a quantity the adversary cannot access in our threat model.
@@
Concrete adversary instantiations. We propose two concrete instantiations of our gradient-crafting adversaries, which we will use to perform privacy auditing on real datasets in the next section (see Appendix C for more precise descriptions):
@@
These two adversaries are designed to be as simple as possible. Other gradient constructions can be used, such as crafting a random gradient following the method proposed by Andrew et al. (2024) (see Figure 8 of the appendix for a comparison with the adversaries we use). However, our choice of adversaries can be further justified by recent findings on privacy backdoors, where the model architecture and parameters are chosen adversarially to leak certain input points. Specifically, Feng and Tramèr (2024) propose a specific construction of architecture and initialization in which a well-chosen input point generates gradients that are nearly concentrated in a single dimension throughout training. This demonstrates that in some cases, our adversaries $\mathcal{A}_{GC}$-R and $\mathcal{A}_{GC}$-S can be instantiated by inserting a canary point.
@@ 5.1 Experimental Setup
Training details. We perform auditing on two datasets: we choose CIFAR10 (Krizhevsky, 2009) as a representative dataset for the state-of-the-art in differentially private training (Tramèr and Boneh, 2021; De et al., 2022), and Housing (Pace and Barry, 1997) to underline the hardness of auditing smaller models in the hidden state. We use predefined hyperparameters for each dataset: on CIFAR10, the batch size is 128, and the learning rate is 0.01, while on Housing, the batch size is 400, and the learning rate is 0.1. Training is done with DP-SGD with no momentum. We use three models: a fully connected neural network (FCNN) for the Housing dataset (Pace and Barry, 1997), a convolutional neural network (CovNet) (LeCun et al., 1989) and a residual network (ResNet) (He et al., 2016) for CIFAR10. A detailed description of the models can be found in Appendix D.
@@
Privacy accounting & auditing. We compute privacy upper and lower bounds for the crafted gradient (for $\mathcal{A}_{GC}$) or canary ($\mathcal{A}_{L}$) as follows. The theoretical privacy upper bound is obtained with the numerical accountant proposed by Gopi et al. (2021). We audit three scenarios for which the accountant gives equivalent privacy guarantees, inserting the crafted gradient or canary at different periodicity $k\in\{1,5,25\}$ and adjusting the time horizon $T$ accordingly. More precisely, when $k=1$, the model is trained for 250 optimization steps, and the crafted gradient or canary is inserted at every step; when $k=5$, the model is trained for 1250 steps, and insertion occurs every $5$ steps, and similarly for $k=25$. We report both the theoretical and empirical epsilons at a fixed $\delta=1e^{-5}$, but we note that we can generate the complete privacy curve $(\varepsilon,\delta(\varepsilon))$, which we present in Figure 11 of the appendix for completeness. For auditing, we rely on the recently proposed scheme based on Gaussian DP (described in Appendix A for completeness) as it allows accurate auditing with a small number of auditing runs (Nasr et al., 2023). Details on training and auditing parameters can be found in Appendix D.

### v3

@@ Abstract
Machine learning models can be trained with formal privacy guarantees via differentially private optimizers such as DP-SGD. In this work, we focus on a threat model where the adversary has access only to the final model, with no visibility into intermediate updates. In the literature, this “hidden state” threat model exhibits a significant gap between the lower bound from empirical privacy auditing and the theoretical upper bound provided by privacy accounting. To challenge this gap, we propose to audit this threat model with adversaries that craft a gradient sequence designed to maximize the privacy loss of the final model without relying on intermediate updates. Our experiments show that this approach consistently outperforms previous attempts at auditing the hidden state model. Furthermore, our results advance the understanding of achievable privacy guarantees within this threat model. Specifically, when the crafted gradient is inserted at every optimization step, we show that concealing the intermediate model updates in DP-SGD does not enhance the privacy guarantees. The situation is more complex when the crafted gradient is not inserted at every step: our auditing lower bound matches the privacy upper bound only for an adversarially-chosen loss landscape and a sufficiently large batch size. This suggests that existing privacy upper bounds can be improved in certain regimes.
@@ 1 Introduction
In this work, we consider a natural relaxation where intermediate models are concealed and only the final model is released. This threat model, often referred to as hidden state, is particularly relevant in practice, encompassing scenarios such as open-sourcing a trained model by publishing its weights. From a theoretical perspective, recent work has demonstrated that the hidden state model can yield significantly improved privacy upper bounds for DP-SGD compared to those derived through standard composition (Ye and Shokri, 2022; Altschuler and Talwar, 2022). This improvement can be attributed to the phenomenon of privacy amplification by iteration (Feldman et al., 2018; Balle et al., 2019): in a nutshell, the privacy of a data point used at earlier stages of the optimization process improves as subsequent steps are performed. However, these results only hold for convex problems, leaving a pivotal question unanswered: Does the privacy of non-convex machine learning problems improve when intermediate models are withheld? Empirical lower bounds obtained through privacy auditing (Nasr et al., 2021; 2023; Steinke et al., 2023) suggest that indeed, concealing intermediary model amplifies privacy, exposing a gap between the empirical lower bound and the theoretical upper bound in the hidden state threat model. However, it remains unclear whether this gap arises from genuine privacy amplification (i.e., a loose upper bound) or from the suboptimality of existing adversaries (i.e., a loose lower bound). This raises another crucial question: How can one design worst-case adversaries when the intermediate models are concealed?
@@
Our contributions. We adapt the gradient-crafting adversarial approach of Nasr et al. (2021) to align with the hidden state model, leveraging the key observation that all privacy accounting techniques for DP-SGD are designed to protect against worst-case gradients. Instead of crafting a data point (canary) that gets added to the training set as in prior attempts to audit the hidden state model (Nasr et al., 2023; Steinke et al., 2023), our adversaries craft a sequence of gradients prior to the execution of the algorithm (i.e., without knowledge of the intermediate models). The crafted gradients are then added to gradients computed on real training points to yield the highest possible privacy loss for the final model. In other words, our adversaries abstract away the canary and directly decide the gradient it would have produced when inserted at a given step of DP-SGD. But how do we pick the gradient sequence leading to the worst-case leakage?
@@
In the scenario where crafted gradients are inserted at every iteration of DP-SGD, we demonstrate that gradient-crafting adversaries which allocate the maximum magnitude permitted by DP-SGD to a single gradient dimension are optimal: they imply privacy lower bounds that match the known upper bounds given by the privacy accountant of Gopi et al. (2021). Therefore, our results reveal that releasing only the final model does not amplify privacy in this regime. In the case of small models, we achieve these tight privacy auditing results by carefully selecting the gradient dimension, whereas for over-parameterized models, the dimension selection can be arbitrary.
@@
When the crafted gradient is not inserted at every step, we find that the above adversaries still outperform canary-crafting adversaries by a significant margin but cannot reach the privacy upper bounds. We show that part of the gap can be attributed to the inability of our adversaries to influence the gradients on real training data in steps where the crafted gradient is not inserted. To address this gap, we design an adversary that crafts both the gradient and the loss landscape. Our results uncover two distinct regimes: (i) when the batch size is large compared to the noise variance, our adversary proves to be optimal, yielding a privacy lower bound that matches the known upper bound and thus demonstrating the absence of privacy amplification; and (ii) when the batch is small relative to the noise variance, we show strong evidence of privacy amplification for non-convex problems, although the effect is qualitatively weaker than in the convex case. Our findings advance the understanding of privacy guarantees in the hidden state model, and lay the groundwork for the design of better privacy accounting techniques for this threat model.
@@ 2.1 Differential Privacy and DP-SGD
The workhorse of private machine learning is the Differentially Privacy Stochastic Gradient Descent (DP-SGD) algorithm (Song et al., 2013; Bassily et al., 2014; Abadi et al., 2016). Let $D$ be the training dataset, $\theta$ the model parameters and consider the standard empirical risk minimization objective $\min_{\theta}\frac{1}{|D|}\sum_{x\in D}\ell(\theta;x)$ where $\ell$ is a loss function differentiable in its first parameter. DP-SGD follows similar steps as standard SGD but ensures differential privacy by (i) bounding the contribution of each data point to the gradient using clipping and (ii) adding Gaussian noise to the clipped gradients. Formally, starting from some initialization $\theta_{0}$, DP-SGD performs $T$ iterative updates of the form:
@@ 2.2 Threat Models for DP-SGD
DP protects against an adversary that observes the output of either $\mathcal{M}(D)$ or $\mathcal{M}(D^{\prime})$ and seeks to predict whether the dataset was $D$ or $D^{\prime}=D\cup\{x\}$, i.e., whether $x$ was included in the input dataset (a.k.a., a membership inference attack) (Homer et al., 2008; Shokri et al., 2017). In this context, the threat model specifies which information is observable/known by the adversary. For DP-SGD, in addition to the final model $\theta_{T}$, threat models typically consider that the adversary has access to (and potentially controls) the model architecture, the loss $\ell$, the initialization $\theta_{0}$, the dataset (up to the presence of $x$) and differ in which internal states of DP-SGD are observable by the adversary. Below, we recall the two threat models relevant to our work.
Standard threat model. Standard privacy accounting techniques analyze DP-SGD as a composition of (potentially subsampled) Gaussian mechanisms (Abadi et al., 2016; Gopi et al., 2021; Doroshenko et al., 2022). Composition allows the adversary to observe all intermediate models $\theta_{1},...,\theta_{T-1}$ (in addition to $\theta_{0}$ and $\theta_{T}$). Previous work has shown that existing privacy accounting techniques are tight in this threat model (Nasr et al., 2021; 2023). However, revealing intermediate models is often unnecessary in practical deployment scenarios (particularly in centralized settings where a single entity conducts the training process) and may degrade the privacy-utility trade-off, thereby motivating the study of the hidden state threat model.
@@ 2.3 Auditing Differential Privacy
Privacy accounting only provides upper bounds on the privacy loss captured by the DP parameters $(\varepsilon,\delta)$, and these bounds are not always tight. Leveraging the relation between DP and the performance of membership inference attacks, privacy auditing aims to produce lower bounds on the DP parameters by instantiating concrete adversaries permitted within a given threat model (Jagielski et al., 2020; Nasr et al., 2021). When these lower bounds match the upper bounds given by privacy accounting, we can conclude that the privacy analysis is tight; when they do not, it is possible to improve the privacy accounting and/or the attacks. A privacy auditing pipeline can be broken down into two components: an adversary and an auditing scheme. The high-level process is shown in Algorithm 1.
@@
     $b_{i}\leftarrow\text{Ber}(1/2)$triangleright draw a random bit
@@ 3 Related Work in Differential Privacy Auditing
Prior attempts at auditing the hidden state model used adversaries employing a loss-based attack that targets a canary obtained by flipping the label of a genuine data point (Nasr et al., 2021; 2023; Steinke et al., 2023; Annamalai and Cristofaro, 2024). The idea is to generate an outlier from an in-distribution data point so that the loss of the model on this canary is high when it is not part of the training set but drops significantly when used during training. Auditing the hidden state model with these adversaries yields privacy lower bounds that exhibit a significant gap compared to privacy accounting upper bounds for the standard threat model (Nasr et al., 2021; 2023; Steinke et al., 2023; Annamalai and Cristofaro, 2024). This gap has two possible explanations: (i) these adversaries are suboptimal and stronger adversaries exist in the hidden state model; and/or (ii) the upper bounds are not tight. Understanding the reasons for this gap, so as to gain better knowledge of the properties of DP-SGD in the hidden state model, is the main motivation for our work. We demonstrate that gradient-crafting adversaries can provide tighter auditing results, thereby proving the validity of hypothesis (i). We show that in certain regimes, our auditing results match the privacy upper bounds, indicating that releasing intermediate models does not increase the privacy loss. However, we also identify regimes where a gap remains, thus providing compelling evidence in support of hypothesis (ii). Specifically, our findings suggest the presence of a privacy amplification by iteration phenomenon in the non-convex setting, albeit weaker compared to what is theoretically established for the convex case (Feldman et al., 2018; Balle et al., 2019; Bok et al., 2024), but less restrictive than the results of Asoodeh and Diaz (2023). The latter require projecting the iterates onto a convex set of bounded diameter or a strong enough decay term on the parameters, which are unrealistic assumptions in practical scenarios involving non-convex and over-parameterized models like the ones considered in our work.
@@ 4 Gradient-Crafting Adversaries for the Hidden State Model
In all threat models, DP-SGD enjoys a simplified interpretation as a sequence of $T$ sum queries with bounded terms, where the $t$-th query corresponds to summing the clipped gradients over the mini-batch $B_{t}$ in Equation 2. From this perspective, privacy accounting techniques for DP-SGD are designed to accommodate any possible gradients with a bounded norm $C$, irrespective of whether they originate from a fixed dataset. This highlights a key limitation of adversaries used so far in the hidden state model: due to the non-convexity of the objective function, any sequence of gradients for the canary $x^{*}$ is likely to be possible, but how to craft a worst-case $x^{*}$ is unclear (and in fact out of scope for privacy auditing). Instead, we allow the adversary to directly craft any gradient sequence, considering that there could be a data point that would generate that sequence. To the best of our knowledge, there is no theoretical restriction on the gradients that neural networks can produce; therefore, we cannot rule out the possibility that some input point could generate an arbitrary sequence of gradients.
@@
  • Saturating the gradient norm: The canary point crafted by prior adversaries is not guaranteed to saturate the gradient clipping threshold throughout training. The adversary’s performance becomes architecture-dependent: for example, a small convolutional neural network has a small canary norm at the start of the training compared to a ResNet, as observed in Figure 8 of the Appendix. Thus, for a tight audit, one must tune the clipping threshold according to the canary gradient norm, which the adversary cannot access in our threat model.
@@
Concrete adversary instantiations. We propose two instantiations of our gradient-crafting adversaries, which we will use to perform privacy auditing in the next section (see ):
@@
These two adversaries are designed to be as simple as possible. Other gradient constructions can be used, such as crafting a random gradient following the method proposed by Andrew et al. (2024) (see Figure 8 of the appendix for a comparison with the adversaries we use). However, recent findings on privacy backdoors can further justify our choice of adversaries, where the model architecture and parameters are chosen adversarially to leak certain input points. Specifically, Feng and Tramèr (2024) propose a specific construction of architecture and initialization in which a well-chosen input point generates gradients nearly concentrated in a single dimension throughout training. This demonstrates that in some cases, our adversaries $\mathcal{A}_{GC}$-R and $\mathcal{A}_{GC}$-S can be instantiated by inserting a canary point.
For more details on our two gradient-crafting adversaries, see Appendix C, and refer to Algorithm 4 for the corresponding auditing procedure, adapted from Algorithm 1.
@@ 5.1 Experimental Setup
+Training details. We perform auditing on two datasets: we choose CIFAR10 (Krizhevsky, 2009) as a representative dataset for the state-of-the-art in differentially private training (Tramèr and Boneh, 2021; De et al., 2022), and Housing (Pace and Barry, 1997) to underline the hardness of auditing smaller models in the hidden state. We use fixed hyperparameters for each dataset: on CIFAR10, the batch size is 128, and the learning rate is 0.01, while on Housing, the batch size is 400, and the learning rate is 0.1. Training is done with DP-SGD with no momentum. We use three models: a fully connected neural network (FCNN) for the Housing dataset (Pace and Barry, 1997), a convolutional neural network (CovNet) (LeCun et al., 1989) and a residual network (ResNet) (He et al., 2016) for CIFAR10. A detailed description of the models can be found in Appendix D.
@@
Privacy accounting & auditing. We compute privacy upper and lower bounds for the crafted gradient (for $\mathcal{A}_{GC}$) or canary ($\mathcal{A}_{L}$) as follows. The theoretical privacy upper bound is given by the accountant of Gopi et al. (2021). We audit three scenarios where the accountant gives equivalent privacy guarantees for the crafted gradient or canary by inserting it at different periodicity $k\in\{1,5,25\}$, accounting only for the steps where the insertion occurs and adjusting the time horizon $T$ accordingly. When $k=1$, the model is trained for $T=250$ steps, and the crafted gradient or canary is inserted at every step; when $k=5$, the model is trained for 1250 steps and insertion occurs every $5$ steps; and similarly for $k=25$. We report both the theoretical and empirical epsilons at a fixed $\delta=1e^{-5}$, but note that we can generate the complete privacy curve $(\varepsilon,\delta(\varepsilon))$, see Figure 11 of the appendix. For auditing, we rely on the recently proposed scheme based on Gaussian DP (described in Appendix A for completeness) as it allows accurate auditing with a small number of auditing runs (Nasr et al., 2023). Details on training and auditing parameters can be found in Appendix D.